in a graph with the points (-5,10) (3,y) and the slope=-(3/8) I need to solve for y

Found 2 solutions by nerdybill, richwmiller:
Answer by nerdybill(7384)

in a graph with the points (-5,10) (3,y) and the slope=-(3/8) I need to solve for y
.
slope, given two points:
m = (y2 - y1)/(x2 - x1)
.
The problem gives us:
m = -3/8
(x1,y1) as (-5,10)
(x2,y2) as (3,y)
.
Plug the given info into:
m = (y2 - y1)/(x2 - x1)
we get
-3/8 = (y - 10)/(3 - (-5))
-3/8 = (y - 10)/(3 + 5)
-3/8 = (y - 10)/8
Multiplying both sides by 8:
-3 = y - 10
7 = y

Answer by richwmiller(17219)

y=mx+b
y=-3/8x+b
plug in x=-5 and y=10
10=-3/8*(-5)+b
solve for b
10=15/8+b
65/8=b
y=-3/8x+65/8
now plug in 3 for x and solve for y
y=-3*3/8+65/8
multiply by 8
8y=9+65
8y=74
y=74/8
y=37/4